# Singular-degenerate parabolic systems with the conormal boundary condition on the upper half space

### v1

@@ 3. $L_{2}$ solvability
In this section, we establish the $L_{2}(\Omega_{T},\mu)$ solvability for systems
@@
Let $\alpha\in(-1,\infty),\kappa\in(0,1],\lambda\geq 0$, and suppose that (2.1) and (2.2) are satisfied. If $u\in\mathcal{H}^{1}_{2}(\Omega_{T},\mu)$ is a weak solution to (3.1)-(3.2) for some $F\in L_{2}(\Omega_{T},\mu)^{d}$ and $f\in L_{2}(\Omega_{T},\mu)$, then $u\in C_{b}(\overline{\mathbb{R}_{T}};L_{2}(\mathbb{R}^{d}_{+},\mu))$ and
$\sup_{t<T}\int_{\mathbb{R}^{d}_{+}}|u(t,\cdot)|^{2}d\mu+\int_{\Omega_{T}}|Du|^{2}d\mu+\lambda\int_{\Omega_{T}}|u|^{2}d\mu\leq N\int_{\Omega_{T}}|F|^{2}d\mu+N\int_{\Omega_{T}}|f|^{2}d\mu$ (3.3)
for some $N=N(d,n,\alpha,\kappa)>0$. Moreover, for any $\lambda>0$, $F\in L_{2}(\Omega_{T},\mu)^{d}$, and $f\in L_{2}(\Omega_{T},\mu)$, there exists a unique weak solution $u\in\mathcal{H}^{1}_{2}(\Omega_{T},\mu)$ to (3.1)-(3.2).
Here $C_{b}(\overline{\mathbb{R}_{T}};L_{2}(\mathbb{R}^{d}_{+},\mu))$ stands for the space of continuous $L_{2}(\mathbb{R}^{d}_{+},\mu)$-valued functions on $\overline{\mathbb{R}_{T}}$. Before giving the proof, we introduce some notation and auxiliary results. We first recall the Steklov averages in the time variable. For $h>0$, the backward Steklov average of a function $u(t,x)$ is the function
@@
We will use the following properties of Steklov averages. If $u\in L_{2}(\Omega_{T},\mu)$, then
$\|u_{h}\|_{L_{2}(\Omega_{T},\mu)}\leq\|u\|_{L_{2}(\Omega_{T},\mu)}\quad\text{for any }h>0,$
@@
$\lim_{h\to 0^{+}}\|u_{h}-u\|_{L_{2}(\Omega_{T},\mu)}=0.$
where $\eta_{1}\in C_{0}^{\infty}(0,1),\eta_{2}\in C_{0}^{\infty}(B_{1}^{\prime})$ are non-negative with unit integrals.
For $T=\infty$, let $\hat{u}(\tau,x):=(2\pi)^{-1/2}\int_{\mathbb{R}}u(t,x)e^{-it\tau}dt$ be the Fourier transform in the time variable of $u\in L_{2}(\mathbb{R}^{d+1}_{+},\mu)$. The Hilbert transform in time $H$ and $(-\partial_{t}^{2})^{1/4}$ operators are defined by
$\widehat{Hu}(\tau,x)=-i\operatorname{sgn}\tau\hat{u}(\tau,x),\quad\widehat{(-\partial_{t}^{2})^{1/4}u}(\tau,x)=|\tau|^{1/2}\hat{u}(\tau,x).$
@@
$H^{1/2,1}_{2}(\mathbb{R}^{d+1}_{+},\mu)=\big\{u\in L_{2}(\mathbb{R}^{d+1}_{+},\mu):(-\partial_{t}^{2})^{1/4}u,Du\in L_{2}(\mathbb{R}^{d+1}_{+},\mu)\big\}$
@@
$\langle u,v\rangle_{H^{1/2,1}_{2}(\mathbb{R}^{d+1}_{+},\mu)}=\int_{\mathbb{R}^{d+1}_{+}}\left[(-\partial_{t}^{2})^{1/4}u\cdot(-\partial_{t}^{2})^{1/4}v+D_{i}u\cdot D_{i}v+u\cdot v\right]d\mu.$
We say that $u\in H^{1/2,1}_{2}(\mathbb{R}^{d+1}_{+},\mu)$ is an $H^{1/2,1}_{2}(\mathbb{R}^{d+1}_{+},\mu)$-solution to (3.4)-(3.2) if
@@
holds for all $v\in H^{1/2,1}_{2}(\mathbb{R}^{d+1}_{+},\mu)$. Clearly, if $g=0$, an $H^{1/2,1}_{2}(\mathbb{R}^{d+1}_{+},\mu)$-solution is also a weak solution to (3.1)-(3.2) in the sense of identity (2.3). We will need the following result on the solvability of (3.4)-(3.2) in $H^{1/2,1}_{2}(\mathbb{R}^{d+1}_{+},\mu)$.
@@
Let $\alpha\in(-1,\infty)$ and suppose that (2.1) and (2.2) are satisfied. Then for any $\lambda>0,F\in L_{2}(\mathbb{R}^{d+1}_{+},\mu)^{d},$ and $f,g\in L_{2}(\mathbb{R}^{d+1}_{+},\mu)$, there exists a unique $H^{1/2,1}_{2}(\mathbb{R}^{d+1}_{+},\mu)$-solution $u$ to (3.4)-(3.2).
@@
for all real-valued $f\in L_{2}(\mathbb{R}^{d+1}_{+},\mu)^{n}$. For the rest of the details, we refer to the proof of [27, Theorem 4.1]. ∎
@@
By taking Steklov averages, for any $h>0$ and $\varphi\in C_{0}^{\infty}((-\infty,T)\times\mathbb{R}^{d})$,
@@
By considering cut-off in time, for any $v\in C_{0}^{\infty}(\mathbb{R}^{d+1})$ and $S\leq T$,
@@
By approximation, (3.5) holds for all $v\in L_{2}(\Omega_{S},\mu)$ with $Dv\in L_{2}(\Omega_{S},\mu)$.
Now take $v=u_{h}$. Then
@@
Therefore, it holds that
@@
$\int_{\Omega_{S}}[A^{ij}D_{j}u\cdot D_{i}u+\lambda C_{0}u\cdot u]d\mu\leq\int_{\Omega_{S}}[F^{i}\cdot D_{i}u+\lambda^{1/2}f\cdot u]d\mu$
for all $S\leq T$. By (2.1), (2.2), and Young’s inequality, we get
@@
From (3.5), (3.6), and (3.7), we get ($\|\cdot\|=\|\cdot\|_{L_{2}(\Omega_{S},\mu)}$)
@@
Hence $\|u_{h_{1}}-u_{h_{2}}\|_{L_{\infty}((-\infty,T);L_{2}(\mathbb{R}^{d}_{+},\mu))}\to 0$ as $h_{1},h_{2}\to 0^{+}$, from which and also (3.8), we have
$u\in C(\overline{\mathbb{R}_{T}};L_{2}(\mathbb{R}^{d}_{+},\mu))$
@@
$\sup_{t<T}\int_{\mathbb{R}_{d}^{+}}|u(t,\cdot)|^{2}d\mu\leq N\|F\|_{L_{2}(\Omega_{T},\mu)}^{2}+\|f\|_{L_{2}(\Omega_{T},\mu)}^{2}.$
Estimate (3.3) is proved. It suffices to prove the existence of solutions for the case $T=\infty$, which follows by Theorem 3.2. ∎
@@ 4. Systems with simple coefficients
Let $\alpha\in(-1,\infty),p\in(1,\infty),\kappa\in(0,1]$, and $\lambda>0$. Suppose that (2.1) and (2.2) are satisfied. Then, for any $F\in L_{p}(\Omega_{T},\mu)^{d}$ and $f\in L_{p}(\Omega_{T},\mu)$, there exists a unique solution $u\in\mathcal{H}^{1}_{p}(\Omega_{T},\mu)$ to (4.1)-(4.2). Moreover,
@@
where $N=N(d,n,\alpha,p,\kappa)>0$.
Throughout this section, it is assumed that (2.1) and (2.2) are satisfied with some $\kappa\in(0,1]$, $\alpha\in(-1,\infty)$, and $\lambda\geq 0$. For compactness of notation, denote
@@ 4.1. Lipschitz estimates for homogeneous systems
$\lim_{x_{d}\to 0^{+}}x_{d}^{\alpha}A^{dj}(x_{d})D_{j}u=0\quad\text{if $\overline{D_{R}^{+}(x_{0})}\cap\partial\mathbb{R}^{d}_{+}\neq\emptyset$}.$ (4.5)
@@
Let $0<r<R,z_{0}=(t_{0},x_{0})\in\overline{\mathbb{R}^{d+1}_{+}}$, and $u\in\mathcal{H}^{1}_{2}(Q_{R}^{+}(z_{0}),\mu)$ be a weak solution to (4.4)-(4.5) in $Q_{R}^{+}(z_{0})$. Then $u_{t}\in L_{2}(Q_{r}^{+}(z_{0}))$ and
@@
where $N=N(d,n,\alpha,\kappa)>0$.
The proof is standard; see, for example, the argument in [11, p. 1082-1084]. For the first inequality, we test the equation by $u\varphi^{2}$, where $\varphi$ is an appropriately chosen smooth cutoff function. For the second inequality, we test the equation by $u_{t}\varphi_{k}^{2}$, where $\varphi_{k}$ are appropriately chosen smooth cutoff functions with expanding supports. We then use the first inequality and an iteration argument. As in the proof of Theorem 3.1, one can use the Steklov averages to justify taking $u\varphi^{2}$ as test functions. By considering finite differences with respect to $t$ and $x^{\prime}$ in equation (4.4) and iterating the estimates in Lemma 4.2, we have
@@
$\int_{Q_{r}^{+}(z_{0})}|\partial_{t}^{I}D_{x^{\prime}}^{J}u|^{2}d\mu\leq\frac{N_{I,J}}{(R-r)^{4I+2J-2}}\int_{Q_{R}^{+}(z_{0})}(|Du|^{2}+\lambda|u|^{2})d\mu,\quad I+J\geq 1,$
$\int_{Q_{r}^{+}(z_{0})}(\lambda^{k}|D\partial_{t}^{I}D_{x^{\prime}}^{J}u|^{2}+\lambda^{k+1}|\partial_{t}^{I}D_{x^{\prime}}^{J}u|^{2})d\mu\leq\frac{N_{I,J,k}}{(R-r)^{4I+2J+2k}}\int_{Q_{R}^{+}(z_{0})}(|Du|^{2}+\lambda|u|^{2})d\mu,\,I,J,k\geq 0.$
Here $N_{I,J}$ and $N_{I,J,k}$ depend only on $d,n,\alpha,\kappa$, and the subscripted indices.
In the following lemma, we prove interior $L_{\infty}$ estimates of solutions. Denote
@@
Let $p\in(1,2],r>0$, $z_{0}\in\mathbb{R}^{d+1}_{+}$ with $x_{0d}\geq 2r$, and let $u\in\mathcal{H}^{1}_{p}(Q_{r}(z_{0}),\mu)$ be a weak solution to (4.4) in $Q_{r}(z_{0})$. Then
$\begin{gathered}\|u\|_{L_{\infty}(Q_{r/2}(z_{0}))}\leq N\left(\fint_{Q_{r}(z_{0})}|u(t,x)|^{p}d\mu(z)\right)^{1/p},\\ \|Du\|_{L_{\infty}(Q_{r/2}(z_{0}))}+\sqrt{\lambda}\|u\|_{L_{\infty}(Q_{r/2}(z_{0})}\leq N\left(\fint_{Q_{r}(z_{0})}(|Du(t,x)|^{p}+\lambda^{p/2}|u(t,x)|^{p})d\mu(z)\right)^{1/p},\end{gathered}$ (4.6)
where $N=N(d,n,\alpha,p,\kappa)>0$.
@@
Since $x_{d}\sim x_{0d}$ in $Q_{r}(z_{0})$, by multiplying the equation with $x_{0d}^{-\alpha}$ and a translation of coordinates, we may assume without loss that $\alpha=0$ and $z_{0}=0$. By scaling, it is sufficient to consider $r=1$. We first consider the case $p=2$. For compactness of notation, let
@@
Fix an even integer $k>(d+1)/2$. By the Sobolev embedding and Lemmas 4.2 and 4.3, for $z\in Q_{1/2}$, it holds that
@@
$\leq N\int_{-1/2}^{1/2}(\|u(\cdot,s)\|^{2}_{W^{k/2,k}_{2}(Q_{1/2}^{\prime})}+\|D_{d}u(\cdot,s)\|^{2}_{W^{k/2,k}_{2}(Q_{1/2}^{\prime})})ds\leq NC_{1}^{2}.$

### v2

@@ 3. $L_{2}$ solvability
In this section, we establish the $L_{2}(\Omega_{T},\mu;\mathbb{R}^{n})$ solvability for systems
@@
Let $\alpha\in(-1,\infty),\kappa\in(0,1],\lambda\geq 0$, and suppose that (2.1) and (2.2) are satisfied. If $u\in\mathcal{H}^{1}_{2}(\Omega_{T},\mu;\mathbb{R}^{n})$ is a weak solution to (3.1)-(3.2) for some $F\in L_{2}(\Omega_{T},\mu;\mathbb{R}^{n})^{d}$ and $f\in L_{2}(\Omega_{T},\mu;\mathbb{R}^{n})$, then $u\in C_{b}(\overline{\mathbb{R}_{T}};L_{2}(\mathbb{R}^{d}_{+},\mu;\mathbb{R}^{n}))$ and
$\sup_{t\in\overline{\mathbb{R}_{T}}}\int_{\mathbb{R}^{d}_{+}}|u(t,\cdot)|^{2}d\mu+\int_{\Omega_{T}}|Du|^{2}d\mu+\lambda\int_{\Omega_{T}}|u|^{2}d\mu\leq N\int_{\Omega_{T}}|F|^{2}d\mu+N\int_{\Omega_{T}}|f|^{2}d\mu$ (3.3)
for some $N=N(\kappa)>0$. Moreover, for any $\lambda>0$, $F\in L_{2}(\Omega_{T},\mu;\mathbb{R}^{n})^{d}$, and $f\in L_{2}(\Omega_{T},\mu;\mathbb{R}^{n})$, there exists a unique weak solution $u\in\mathcal{H}^{1}_{2}(\Omega_{T},\mu;\mathbb{R}^{n})$ to (3.1)-(3.2).
Here $C_{b}(\overline{\mathbb{R}_{T}};L_{2}(\mathbb{R}^{d}_{+},\mu;\mathbb{R}^{n}))$ stands for the space of continuous $L_{2}(\mathbb{R}^{d}_{+},\mu)$-valued functions on $\overline{\mathbb{R}_{T}}$. Before giving the proof, we introduce some notation and auxiliary results. We first recall the Steklov averages in the time variable. For $h>0$, the backward Steklov average of a function $u(t,x)$ is the function
@@
We will use the following properties of Steklov averages. If $u\in L_{2}(\Omega_{T},\mu;\mathbb{R}^{n})$, then
$\|u_{h}\|_{L_{2}(\Omega_{T},\mu;\mathbb{R}^{n})}\leq\|u\|_{L_{2}(\Omega_{T},\mu;\mathbb{R}^{n})}\quad\text{for any }h>0,$
@@
$\lim_{h\to 0^{+}}\|u_{h}-u\|_{L_{2}(\Omega_{T},\mu;\mathbb{R}^{n})}=0.$
For $T=\infty$, let $\hat{u}(\tau,x):=(2\pi)^{-1/2}\int_{\mathbb{R}}u(t,x)e^{-it\tau}dt$ be the Fourier transform in the time variable of $u\in L_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{C}^{n})\cap L_{1}(\mathbb{R};L_{2}(\mathbb{R}^{d}_{+},\mu;\mathbb{C}^{n}))$. By density and Plancherel’s theorem, the Fourier transform in time extends to an isometric isomorphism $\mathcal{F}_{t}:L_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{C}^{n})\to L_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{C}^{n})$. The Hilbert transform in time $H$ is the isometric isomorphism $L_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{C}^{n})\to L_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{C}^{n})$ defined by
$\widehat{Hu}(\tau,x)=-i\operatorname{sgn}\tau\;\hat{u}(\tau,x).$
For functions $u\in L_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{C}^{n})$ with $|\tau|^{1/2}\hat{u}\in L_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{C}^{n})$, the $(-\partial_{t}^{2})^{1/4}$ operator is defined by
$(-\partial_{t}^{2})^{1/4}u=\mathcal{F}_{t}^{-1}(|\tau|^{1/2}\hat{u})\in L_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{C}^{n}).$
A straightforward verification shows that if $u\in L_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{R}^{n})$ is $\mathbb{R}^{n}$-valued, then $Hu\in L_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{R}^{n})$ is also $\mathbb{R}^{n}$-valued. Moreover, if $u\in L_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{R}^{n})$ is $\mathbb{R}^{n}$-valued and $|\tau|^{1/2}\hat{u}\in L_{2}(\mathbb{R}^{d+1}_{+},\mu)$, then $(-\partial_{t}^{2})^{1/4}u\in L_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{R}^{n})$ is also $\mathbb{R}^{n}$-valued.
@@
$H^{1/2,1}_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{R}^{n})=\big\{u\in L_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{R}^{n}):(-\partial_{t}^{2})^{1/4}u,D_{1}u,\dots,D_{d}u\in L_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{R}^{n})\big\}$
@@
$\langle u,v\rangle_{H^{1/2,1}_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{R}^{n})}=\int_{\mathbb{R}^{d+1}_{+}}\left[(-\partial_{t}^{2})^{1/4}u\cdot(-\partial_{t}^{2})^{1/4}v+D_{i}u\cdot D_{i}v+u\cdot v\right]d\mu.$
We say that $u\in H^{1/2,1}_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{R}^{n})$ is an $H^{1/2,1}_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{R}^{n})$-solution to (3.4)-(3.2) if
@@
holds for all $v\in H^{1/2,1}_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{R}^{n})$. Clearly, if $g=0$, an $H^{1/2,1}_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{R}^{n})$-solution is also a weak solution to (3.1)-(3.2) in the sense of identity (2.3). We will need the following result on the solvability of (3.4)-(3.2) in $H^{1/2,1}_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{R}^{n})$.
@@
Let $\alpha\in(-1,\infty)$ and suppose that (2.1) and (2.2) are satisfied. Then for any $\lambda>0,F\in L_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{R}^{n})^{d},$ and $f,g\in L_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{R}^{n})$, there exists a unique $H^{1/2,1}_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{R}^{n})$-solution $u$ to (3.4)-(3.2).
@@
for all $f\in L_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{R}^{n})$. For the rest of the details, we refer to the proof of [31, Theorem 4.1]. ∎
@@
By taking Steklov averages, for any $h>0$ and $\varphi\in C_{0}^{\infty}((-\infty,T)\times\mathbb{R}^{d};\mathbb{R}^{n})$,
@@
By considering cut-off in time, for any $v\in C_{0}^{\infty}(\mathbb{R}^{d+1};\mathbb{R}^{n})$ and $S\leq T$,
@@
By approximation, (3.5) holds for all $v\in L_{2}(\Omega_{S},\mu;\mathbb{R}^{n})$ with $Dv\in L_{2}(\Omega_{S},\mu)$.
Now take $v=u_{h}$. Then for $S\in\overline{\mathbb{R}_{T}}$ holds
@@
If $S=T=\infty$, then
$\int_{\mathbb{R}^{d+1}_{+}}a_{0}(x)\partial_{t}u_{h}\cdot u_{h}d\mu=0.$
Therefore, for all $S\leq T$ holds
@@
$\int_{\Omega_{S}}[A^{ij}D_{j}u\cdot D_{i}u+\lambda C_{0}u\cdot u]d\mu\leq\int_{\Omega_{S}}[F^{i}\cdot D_{i}u+\lambda^{1/2}f\cdot u]d\mu.$
By (2.1), (2.2), and Young’s inequality, we get
@@
Here and throughout the proof $N=N(\kappa)>0$. From (3.5), (3.6), and (3.7), for all $S\in\overline{\mathbb{R}_{T}}$ it holds that ($\|\cdot\|=\|\cdot\|_{L_{2}(\Omega_{S},\mu)}$)
@@
Hence $\|u_{h_{1}}-u_{h_{2}}\|_{C_{b}(\overline{\mathbb{R}_{T}};L_{2}(\mathbb{R}^{d}_{+},\mu))}\to 0$ as $h_{1},h_{2}\to 0^{+}$, from which and also (3.8), we have
$u\in C_{b}(\overline{\mathbb{R}_{T}};L_{2}(\mathbb{R}^{d}_{+},\mu))$
@@
$\sup_{t\in\overline{\mathbb{R}_{T}}}\int_{\mathbb{R}_{d}^{+}}|u(t,\cdot)|^{2}d\mu\leq N\|F\|_{L_{2}(\Omega_{T},\mu)}^{2}+N\|f\|_{L_{2}(\Omega_{T},\mu)}^{2}.$
Estimate (3.3) is proved. By extension, it suffices to prove the existence of solutions for the case $T=\infty$, which follows by Theorem 3.2. ∎
@@ 4. Systems with simple coefficients
Let $\alpha\in(-1,\infty),p\in(1,\infty),\kappa\in(0,1]$, and $\lambda>0$. Suppose that (2.1) and (2.2) are satisfied. Then, for any $F\in L_{p}(\Omega_{T},\mu;\mathbb{R}^{n})^{d}$ and $f\in L_{p}(\Omega_{T},\mu;\mathbb{R}^{n})$, there exists a unique solution $u\in\mathcal{H}^{1}_{p}(\Omega_{T},\mu;\mathbb{R}^{n})$ to (4.1)-(4.2). Moreover,
@@
where $N=N(d,\alpha,p,\kappa)>0$.
Throughout this section, it is assumed that $\alpha\in(-1,\infty)$, $\lambda\geq 0$, and that (2.1) and (2.2) are satisfied with some $\kappa\in(0,1]$. For compactness of notation, denote
@@ 4.1. Lipschitz estimates for homogeneous systems
$\lim_{x_{d}\to 0^{+}}x_{d}^{\alpha}A^{dj}(x_{d})D_{j}u=0\quad\text{on $Q_{R}^{\prime}(z_{0}^{\prime})$ if $x_{0d}\leq R$}.$ (4.5)
@@
Let $0<r<R,z_{0}=(t_{0},x_{0})\in\overline{\mathbb{R}^{d+1}_{+}}$, and $u\in\mathcal{H}^{1}_{2}(Q_{R}^{+}(z_{0}),\mu;\mathbb{R}^{n})$ be a weak solution to (4.4)-(4.5) in $Q_{R}^{+}(z_{0})$. Then $u_{t}\in L_{2}(Q_{r}^{+}(z_{0}))$ and
@@
where $N=N(\kappa)>0$.
The proof is standard; see, for example, the argument in [12, p. 1082-1084]. For the first inequality, we test the equation by $u\varphi^{2}$, where $\varphi$ is an appropriately chosen smooth cutoff function. As in the proof of Theorem 3.1, one can use the Steklov averages to justify taking $u\varphi^{2}$ as a test function. For the second inequality, note that Steklov averages $u_{h}$ are solutions to (4.4)-(4.5) on $(t_{0}-R^{2}+h,t_{0})\times D_{R}^{+}(x_{0})$. We then test the equation for $u_{h}$ by $(\partial_{t}u_{h})\varphi_{k}^{2}$, where $\varphi_{k}$ are appropriately chosen smooth cutoff functions with expanding supports, and use an iteration argument combined with the first inequality. This proves the second inequality with $u_{h}$ instead of $u$ on the left hand side. We then pass to a limit along a subsequence as $h\to 0^{+}$. Note that the second inequality holds if the coefficients are merely measurable and independent of $t$, and for the first inequality, we used only that $a_{0}$ does not depend on $t$.
When the coefficients are independent of $t$ and $x^{\prime}$, we can iterate the estimates in Lemma 4.2 to show that all partial derivatives of $u$ in $t,x^{\prime}$ exist and are again solutions to (4.4)-(4.5). For $I,J\geq 0$, we denote
$|\partial_{t}^{I}D_{x^{\prime}}^{J}u|^{2}=\sum_{j_{1},\dots,j_{J}=1}^{d-1}|\partial_{t}^{I}D_{j_{1},\dots,j_{J}}u|^{2},\quad|D\partial_{t}^{I}D_{x^{\prime}}^{J}u|^{2}=\sum_{j=1}^{d}\sum_{j_{1},\dots,j_{J}=1}^{d-1}|\partial_{t}^{I}D_{j,j_{1},\dots,j_{J}}u|^{2}.$
where $\eta_{1}\in C_{0}^{\infty}(0,1),\eta_{2}\in C_{0}^{\infty}(B_{1}^{\prime})$ are fixed non-negative functions with unit integrals.
@@
$\int_{Q_{r}^{+}(z_{0})}|\partial_{t}^{I}D_{x^{\prime}}^{J}u|^{2}d\mu\leq\frac{N_{\kappa,I,J}}{(R-r)^{4I+2J-2}}\int_{Q_{R}^{+}(z_{0})}(|Du|^{2}+\lambda|u|^{2})d\mu,\quad I+J\geq 1,$
$\int_{Q_{r}^{+}(z_{0})}(\lambda^{k}|D\partial_{t}^{I}D_{x^{\prime}}^{J}u|^{2}+\lambda^{k+1}|\partial_{t}^{I}D_{x^{\prime}}^{J}u|^{2})d\mu\leq\frac{N_{\kappa,I,J,k}}{(R-r)^{4I+2J+2k}}\int_{Q_{R}^{+}(z_{0})}(|Du|^{2}+\lambda|u|^{2})d\mu,\,I,J,k\geq 0.$
These inequalities are first proved with $u^{(\varepsilon)}$ instead of $u$ in the left hand sides for all small enough $\varepsilon$ by iterating the estimates in Lemma 4.2. We then pass to a limit along a subsequence as $\varepsilon\to 0^{+}$.
Before we proceed further, we introduce some notation and auxiliary inequalities. Denote
@@
When estimating $Du$, from the structure of the equations, it is often convenient to first estimate $\mathcal{U}$ and $D_{x^{\prime}}u$ separately. From (2.1), we have
$\begin{gathered}\kappa|D_{d}u|\leq|A^{dd}D_{d}u|=|\mathcal{U}-\sum_{j=1}^{d-1}A^{dj}D_{j}u|\leq|\mathcal{U}|+\kappa^{-1}|D_{x^{\prime}}u|,\\ |\mathcal{U}|\leq\kappa^{-1}|Du|.\end{gathered}$ (4.6)
For $p\in[1,\infty)$, even $k\geq 0$, $r>0$, and $u:Q_{r}^{\prime}\to\mathbb{R}^{n}$, we set
$\|u\|^{p}_{W^{k/2,k}_{p}(Q_{r}^{\prime})}:=\int_{Q_{r}^{\prime}}\sum_{2I+J\leq k}|\partial_{t}^{I}D_{x^{\prime}}^{J}u|^{p}dx^{\prime}dt.$
When $k>(d+1)/p$, we have the Sobolev embedding $W^{k/2,k}_{p}(Q_{r}^{\prime};\mathbb{R}^{n})\hookrightarrow L_{\infty}(Q_{r}^{\prime};\mathbb{R}^{n})$ with
$\sup_{Q_{r}^{\prime}}|u|^{p}\leq N_{d,p}\fint_{Q_{r}^{\prime}}\sum_{2I+J\leq k}|r^{2I+J}\partial_{t}^{I}D_{x^{\prime}}^{J}u|^{p}dx^{\prime}dt.$
For scalar-valued functions the above inequality is well-known. For $\mathbb{R}^{n}$-valued functions $u$, one can apply the inequality to scalar functions $u(t,x^{\prime})\cdot\xi$, where $\xi\in\mathbb{R}^{n}$ is fixed, and take $\sup$ over all $\xi$ with $|\xi|=1$. Hence, the constant in the above Sobolev inequality does not depend on $n$.
Lastly, denote $d\mu_{\alpha}=s^{\alpha}ds$ on $\mathbb{R}_{+}$. A direct verification shows that whenever $0\leq a_{1}\leq a_{2}<b_{2}\leq b_{1}$, we have
$\frac{\mu_{\alpha}((a_{1},b_{1}))}{\mu_{\alpha}((a_{2},b_{2}))}=\frac{b_{1}^{1+\alpha}-a_{1}^{1+\alpha}}{b_{2}^{1+\alpha}-a_{2}^{1+\alpha}}\leq N_{\alpha}\left(\frac{b_{1}-a_{1}}{b_{2}-a_{2}}\right)^{1+\alpha_{+}},$ (4.7)
where $N_{\alpha}=\max((1+\alpha)^{-1},1)$ is sharp. In particular, the measure $\mu_{\alpha}$ is doubling.
In the following lemma, we prove interior reverse Hölder type $L_{\infty}$ estimates of solutions.
@@
Let $p\in(1,2],r>0$, $z_{0}\in\mathbb{R}^{d+1}_{+}$ with $x_{0d}\geq 2r$, and let $u\in\mathcal{H}^{1}_{p}(Q_{r}(z_{0}),\mu;\mathbb{R}^{n})$ be a weak solution to (4.4) in $Q_{r}(z_{0})$. Then for any $\tau\in(0,1)$ it holds that
$\begin{gathered}\|u\|_{L_{\infty}(Q_{\tau r}(z_{0}))}\leq N(1-\tau)^{-\frac{d+2}{p}}\left(\fint_{Q_{r}(z_{0})}|u|^{p}d\mu\right)^{1/p},\\ \|Du\|_{L_{\infty}(Q_{\tau r}(z_{0}))}+\sqrt{\lambda}\|u\|_{L_{\infty}(Q_{\tau r}(z_{0}))}\leq N(1-\tau)^{-\frac{d+2}{p}}\left(\fint_{Q_{r}(z_{0})}(|Du|^{p}+\lambda^{p/2}|u|^{p})d\mu\right)^{1/p},\end{gathered}$ (4.8)
where $N=N(d,\alpha,p,\kappa)>0$.
@@
Since $1/2\leq x_{d}/x_{0d}\leq 3/2$ in $Q_{r}(z_{0})$, by multiplying the equation with $x_{0d}^{-\alpha}$ and a translation of coordinates, we may assume without loss that $\alpha=0$ and $z_{0}=0$. Moreover, it suffices to consider the case when $\tau\in[1/2,1)$.
We first consider the case $p=2$ and $\tau=1/2$. By scaling, it is sufficient to consider $r=1$. For compactness of notation, let
@@
Fix the smallest even integer $k$ with $k>(d+1)/2$. By the Sobolev embedding and Lemmas 4.2 and 4.3, for $z\in Q_{1/2}$, it holds that
@@
$\leq N_{d}\int_{-1/2}^{1/2}(\|u(\cdot,s)\|^{2}_{W^{k/2,k}_{2}(Q_{1/2}^{\prime})}+\|D_{d}u(\cdot,s)\|^{2}_{W^{k/2,k}_{2}(Q_{1/2}^{\prime})})ds\leq N_{d,\kappa}C_{1}^{2}.$